# RTGS: Enabling Real-Time Gaussian Splatting on Mobile Devices Using Efficiency-Guided Pruning and Foveated Rendering

### v2

@@ 5. Experimental Setup
We use Mini-Splatting-D (fang2024mini, ), the current-best in quality, as the baseline dense PBNR model. The $L_{1}$ model is obtained from the dense model through pruning and scale decay as described in Sec. 3.4, with an iteration budget of 50,000, followed by another 5,000 iterations of fine-tuning with HVSQ loss. The three lower-quality models are obtained from their immediately higher-quality model as described in Sec. 4.3, with a 7,500 iteration budget.
@@
We design three variants of our method, namely RTGS-H, RTGS-M, and RTGS-L, with decreasing rendering quality. The $L_{1}$ model in the three variants is pruned to have a PSNR of 99%, 98%, and 97% of that of the dense model. The total model size of the three variants is 16%, 12%, and 10%, respectively, of that of the dense model.
@@
We evaluate three real-world datasets: Mip-Nerf360 (barron2022mip, ), Tanks & Temple (Knapitsch2017, ), and DeepBlending (hedman2018deep, ), which amounts to 13 traces in total. The camera poses in the datasets are usually very sparsely populated, which is not representative of the continuous rendering scenario (e.g., VR). We interpolate between the poses in the dataset to create smooth trajectories, producing approximately 1,440 poses for each trace, corresponding to a 16-second video at 90 FPS.
@@
We also compare with two FR methods applied to PBNR. Both methods use the same quality-region division as in our method. The first one is SMFR (Single-Model FR), which uses a single dense PBNR model, which is the $L_{1}$ model in RTGS-H, and randomly samples the points when rendering lower-quality regions. It is effectively a strict subsetting version of our model without selective multi-versioning. The second one is MMFR (Multi-Model FR) (deng2022fov, ), whose $L_{1}$ model is the same as that of RTGS-H and whose higher-level models are pruned from its $L_{1}$ model separately (without subsetting). The number of points used in each level in both methods matches that used in our method.
@@
To assess the subjective rendering quality of our method, we perform a user study; the procedure is approved by our Internal Review Board (IRB). We recruited 12 participants (8 males and 4 females between 20 and 30 years old), all with normal or corrected-to-normal vision. We select four scenes from the three datasets: bicycle, room, drjohnson, and truck, which vary in both content and complexity. We then render the scenes using our RTGS-H and Mini-Splatting-D, which, recall, is the current-best in rendering quality.
@@
We use the classic Two-Interval Forced Choice (2IFC) procedure (perez2019pairwise, ; chen2024pea, ), which is commonly used in psychophysics to compare the quality of two videos. For each trace, we display its rendering by the two methods on the headset in a random order to each participant, with a 5-second rest interval in-between. Each participant is then asked to pick which of the two versions they prefer. Each trace is repeated eight times, and the repetitions across traces are also randomized. The entire experiment lasts about one hour for each participant.
@@ 6. Evaluation
Fig. 9. Performance and objective rendering quality (PSNR) comparison across the five baselines and the three RTGS variants on the mobile Volta GPU. 3DGS and Mini-Splatting-D are dense models, and the other three baselines are pruned models.
@@
6.2. Mobile GPU Results
We now show the performance results on the mobile Volta GPU on Nvidia Jetson AGX Xavier (xaviersoc, ), a representative mobile device for use-cases such as VR. Fig. 10 shows the results. We execute each model five times for each camera pose in each scene in all the datasets, and report the average FPS.
To put the performance results in context, we also compare the PSNR metric of our variants with the baselines. PSNR is a good objective quality measure of the region under the user’s gaze and is universally reported in prior work. Using an objective metric also allows us to scale up the study to more traces. The SSIM results are similar and omitted.
Our three variants Pareto-dominate the baselines. Our slowest variant RTGS-H is 1.9$\times$ faster than the fastest baseline while being 0.1 dB higher than the best-quality baseline. The fastest variant RTGS-L is 7.9$\times$ faster than 3DGS, and can be up to 19.8$\times$ on the largest bicycle trace.
@@ 6.2. Mobile GPU Results
We now ablate the contribution of various performance-enhancing techniques. Fig. 10 shows the FPS (left $y$-axis) and PSNR (right $y$-axis) under: 1) the dense Mini-Splatting-D model, 2) RTGS with only scale decay (SD; Sec. 3.3), 3) RTGS with SD and CE-based pruning (Sec. 3.2), and 4) RTGS with SD, CE, and FR (Sec. 4). We use the RTGS-H model and obtain the FPS/PSNR results on Xavier averaged over all traces.
@@
RTGS-H 102.2 (0.81$\times$) 171.8 (1.06$\times$) 2.12 2.10 2.09 2.08
@@ 6.3. Comparison with Other FR Methods
RTGS also out-performs the two FR baselines. Tbl. 1 compares the FPS (on the mobile Volta GPU), storage requirement, and the HVSQ metric across different quality regions/layers. The results are averaged across all the datasets.
@@
Our method selectively multi-versioned four trainable parameters out of about 60 (Sec. 4.2 and D in Fig. 7), so the additional storage requirement is small (about 6%). Note that RTGS-H already reduces the dense model size to 16% as shown in Sec. 5.
@@ 7. Related Work
The graphics community has long exploited FR for real-time rendering (patney2016towards, ; guenter2012foveated, ; Chen2022InstantReality, ; Konrad2020GazeContingent, ; Kaplanyan2019Deepfovea, ; Krajancich2020Optimizing, ; Chakravarthula2021GazeContingent, ; Sun2017PerceptuallyGuided, ). Particularly relevant to our work, researchers have started applying FR to neural rendering (deng2022fov, ; rolff2023vrs, ; rolff2023interactive, ), such as Fov-NeRF (deng2022fov, ). Our work differs from them in two key ways. First, these methods exclusively focus on NeRF, whereas we focus on PBNR, which is shown to be fundamentally more efficient that NeRF. Second, some (deng2022fov, ) use the multi-model approach, similar to our MMFR baseline, which we out-perform (Sec. 6.3). Our FR method uses subsetting with selectively multi-versioning, addressing the performance overhead of evaluating multiple models (Sec. 4.1).
While conventional FR uses heuristics (e.g., blurring) to guide quality relaxation, recently researchers have investigated more principled ways to model human perception for quality relaxation (FreemanSimoncelli2011, ; walton2021beyond, ). This work leverages such theoretical work and integrates it into the training framework to demonstrate its practical utility.
Prior work has also leveraged eccentricity-dependent color perception of HVS to reduce display power and improve framebuffer compresison in VR (duinkharjav2022color, ; ujjainkar2024exploiting, ; chen2023imperceptible, ; chen2024pea, ), which is orthogonal to this work. A number of early VR video streaming frameworks assume that a user’s gaze is always at the center of the display and reduces the streaming resolution of peripheral regions to save bandwidth. We refer interested readers to EVR (leng2019energy, ) for one such example, which also includes a summary of work in that area.
@@
Almost all existing work optimizing PBNR focuses on pruning, based on the observation that a considerable amount of points can be pruned without impacting the rendering quality. They usually do so by, e.g., explicitly training a mask to remove points (lee2024compact, ) or sorting points by their numerical contribution to pixel colors followed by removing low-contributing points (fan2023lightgaussian, ; girish2023eagles, ; niemeyer2024radsplat, ; fang2024mini, ). People have also investigated non-pruning methods, such as vector quantization (fan2023lightgaussian, ) and distillation (lee2024compact, ) techniques, to compress PBNR models.

### v3

@@ 4.3. HVS-Guided Training
5. Hardware Support
Complementing pruning and FR techniques, we propose an accelerator to further improve the performance. We first provide an overview (Sec. 5.1) and discuss how the hardware addresses the low hardware utilization issue in PBNR, which is exacerbated by FR (Sec. 5.2).
5.1. Overview
Our architecture is built on top of GSCore (Lee et al., 2024a), a recent PBNR accelerator without FR. The basic architecture is designed to support the three PBNR stages discussed in Sec. 2.1. The three stages are pipelined across different tiles in a frame. Fig. 8 shows the pipelined architecture, with the colored components denoting our augmentations. The top panel in Fig. 10 illustrates the pipelining process (omitting the Projection stage for simplicity).
Supporting FR.
We propose two sets of hardware augmentations to support the two new stages in FR (green in panel E). The augmentations are colored in yellow in Fig. 8. First, during the Projection stage we must filter out points that are not used to render a particular quality level. To that end, we augment the Projection hardware to include a filtering unit, which compares the quality bound $m$ of a tile with the current quality level $t$ of that tile, and pushes the tile to the output buffer only if $t>m$.
Second, FR requires blending across quality levels (Sec. 4.1). We augment the Rasterization stage with a blending unit, which takes the two colors of a pixel (rendered by two models corresponding to the quality levels) and performs an interpolation. A small buffer is also needed to temporarily store pixels before blending.
(a) Heatmap showing the number of intersections per tile in bicycle.
(b) Boxplot of the intersection distribution in five traces (clipped at 1,500).
Fig. 9. Workload imbalance, quantified by the number of intersections per tile, on the Mip-NeRF 360 dataset (Barron et al., 2022). The top and bottom notches in the boxplot represent data points that are 1.5 Interquartile Range (IQR) above the third quartile and below the first quartile, respectively.
5.2. Addressing Load Imbalance
The Issue.
With the augmentations above, the hardware can functionally support FR, but faces low hardware utilization. This is because the amount of work each tile requires is severely imbalanced (recall a tile, like an instruction in a conventional processor, is a basic unit for pipelining; Fig. 10).
The amount of work a tile involves can be quantified by the number of tile-ellipse intersections. To quantify this imbalance, Fig. 9a is a heatmap plotting the number of intersections each tile (16$\times$16 pixels) has when rendering an image in the Mip-NeRF 360 dataset using our FR model (with four quality levels). The amount of intersections can vary by over three orders of magnitude. In particular, most of the intersections are concentrated at the center, because the peripheral tiles are rendered using pruned models, which have fewer points. Fig. 9b is the boxplot showing large the distribution of intersections across all traces in the Mip-NeRF 360 dataset (Barron et al., 2022); the imbalance issue is universal.
Load imbalance leads to frequent pipeline stalls. This is illustrated by the top panel in Fig. 10, which shows the pipeline dynamics when pipelining Sorting and Rasterization across four imbalanced tiles. We propose two techniques to address this issue: tile merging and incremental pipelining. The corresponding hardware components are blue-colored in Fig. 8.
Tile Merging.
One way to balance the workload across tiles is to merge tiles that have few intersections. This is dealt with by the Tile Merge Unit (TMU) in the Sorting stage. Conceptually, the TMU merges incoming tiles into a single tile if the cumulative intersections is below a threshold $\beta$. The second panel in Fig. 10 shows an example where the second and third tiles are merged, which reduces pipeline stalls and improves performance.
The TMU has two stages. The first stage processes a Gaussian point by incrementing its counter associated with a specific tile ID, storing the result in a temporal buffer. When the accumulation for one tile is complete, its total count is forwarded to the second stage, which continuously aggregates incoming tiles, accumulates the intersection counts, and compares the cumulative count against $\beta$. If this threshold is exceeded, a merged-tile is formed, where each constituting tile is augmented with a merged-tile ID, which is sent along with the native tile ID to the sorting unit.
Fig. 10. The baseline pipeline faces frequent stalls when the workload is imbalanced across tiles. Tile Merging (TM) and Incremental Pipelining (IP) mitigate the workload imbalanced issue and improves the pipelining efficiency. With IP, when the first sub-tile $S_{11}$ in the $S_{1}$ tile is available by the Sorting stage it can be processed by the Rasterization stage.
Incremental Pipelining.
While Tile Merging reduces the workload imbalances, it does not completely eliminate pipeline stalls, because it is unlikely all the merged tiles have exactly the same number of intersections.
To further enhance pipelining efficiency, we propose to incrementally pipeline data between adjacent stages. In our baseline PBNR accelerator, a double buffer is placed between two adjacent stages; the consumer stage does not start until the producer stage has finished processing the entire tile. Our idea is to break the workload of a tile into smaller sub-tiles so that the consumer stage can start working on available sub-tiles before the entire tile is ready from the producer stage. This works because the workload of each pixel is independent. This is akin to classic superpipelining in processor design (Mirapuri et al., 1992; Shen and Lipasti, 2013). The last panel in Fig. 10 illustrates the benefit of incremental pipelining.
To support such an incremental computation, we replace the double buffers between stages with line buffers (LB) (Hegarty et al., 2014; Redgrave et al., 2018; Hennessy and Patterson, 2019; Ujjainkar et al., 2023), which, under the surface, is a set of small SRAMs, each of which buffers one row of pixels. Assume a 16$\times$16 tile size, once 16 rows are produced in the LB, the consumer can starting using them. The line buffer can be small, since it has to buffer only sub-tiles rather than entire tile.
@@ 6. Experimental Setup
We use Mini-Splatting-D (Fang and Wang, 2024), the current-best in quality, as the baseline dense PBNR model. The $L_{1}$ model is obtained from the dense model through pruning and scale decay as described in Sec. 3.4, with an iteration budget of 50,000, followed by another 5,000 iterations of fine-tuning with HVSQ loss. The three lower-quality models are obtained from their immediately higher-quality model as described in Sec. 4.3, with a 7,500 iteration budget. Our training time is roughly three times as much as that of the original, dense model. Much of the slow down is because the HVS loss is calculated using an open-source Python implementation (hvs, 2022), which could be accelerated with a more efficient implementation in, e.g., CUDA.
@@
We design three variants of our method, namely MetaSapiens-H, MetaSapiens-M, and MetaSapiens-L, with decreasing rendering quality. The $L_{1}$ model in the three variants is pruned to have a PSNR of 99%, 98%, and 97% of that of the dense model. The total model size of the three variants is 16%, 12%, and 10%, respectively, of that of the dense model.
@@
We evaluate three real-world datasets: Mip-Nerf360 (Barron et al., 2022), Tanks & Temple (Knapitsch et al., 2017), and DeepBlending (Hedman et al., 2018), which amounts to 13 traces in total. The camera poses in the datasets are usually very sparsely populated, which is not representative of the continuous rendering scenario (e.g., VR). We interpolate between the poses in the dataset to create smooth trajectories, producing approximately 1,440 poses for each trace, corresponding to a 16-second video at 90 FPS.
Hardware Implementation.
We develop a RTL implementation of the accelerator, where the basic pipeline (the uncolored in Fig. 8) is similar to GScore (Lee et al., 2024a), with the resource allocation adjusted for a more balanced pipeline for our workloads (8 Culling and Conversion Units, a single Hierarchical Sorting Unit, and a 16$\times$16 Volume Rendering Core array). Our RTL design is implemented via Synposys synthesis and Cadence layout tools in TSMC 16nm FinFET technology. Each line buffer has a capacity of 1 KB, and the double buffer before the sorting unit is 64 KB. The SRAMs are generated by an Arm compiler. The DRAM is modeled after four channels of Micron 16 Gb LPDDR3-1600 memory (mic, 2014).
Overall, we have an area of 2.73 mm2. The volume Rendering Core takes 63% of the total area, other stages occupy the rest 30 %; the SRAMs comprise 7% of the total area. Our area is larger than that of GScore (1.45 mm2), whose area is scaled to 16nm using the DeepScaleTool (Sarangi and Baas, 2021). We will show in Sec. 7.5 that we outperform GScore even under the same area when the latter is scaled up.
@@
  • Dense PBNR models: 3DGS (Kerbl et al., 2023), which is the earliest PBNR model, and Mini-Splatting-D (Fang and Wang, 2024), Mip-Spla- tting (Yu et al., 2024), StopThePop (Radl et al., 2024), which are state-of-the-art work that improve upon 3DGS.
  • Pruned PBNR models: LightGS (Fan et al., 2024), CompactGS (Lee et al., 2024b), and Mini-Splatting (Fang and Wang, 2024). The first two are pruned from 3DGS and the last one is from Mini-Splatting-D.
We also compare with two FR methods applied to PBNR. Both methods use the same quality-region division as in our method. The first one is SMFR (Single-Model FR), which uses a single dense PBNR model, which is the $L_{1}$ model in MetaSapiens-H, and randomly samples the points when rendering lower-quality regions. It is effectively a strict subsetting version of our model without selective multi-versioning. The second one is MMFR (Multi-Model FR) (Deng et al., 2022), whose $L_{1}$ model is the same as that of MetaSapiens-H and whose higher-level models are pruned from its $L_{1}$ model separately (without subsetting). The number of points used in each level in both methods matches that used in our method.
@@
To assess the subjective rendering quality of our method, we perform a user study; the procedure is approved by our Internal Review Board (IRB). We recruited 12 participants (8 males and 4 females between 20 and 30 years old), all with normal or corrected-to-normal vision. The scale of our user study is comparable with other research in the field (Deng et al., 2022; Rolff et al., 2023b; Shi et al., 2024; Ye et al., 2022). We select four scenes from the three datasets: bicycle, room, drjohnson, and truck, which vary in both content and complexity. We then render the scenes using our MetaSapiens-H and Mini-Splatting-D, which, recall, is the current-best in rendering quality.
@@
We use the classic Two-Interval Forced Choice (2IFC) psychophysical procedure (Perez-Ortiz et al., 2019; Chen et al., 2024), which is commonly used for evaluating the subjective quality of foveated rendering (Guenter et al., 2012; Walton et al., 2021; Deng et al., 2022; Rolff et al., 2023b; Shi et al., 2024; Zhang et al., 2024; Ye et al., 2022; Wang et al., 2024) For each trace, we display its rendering by the two methods on the headset in a random order to each participant, with a 5-second rest interval in-between. Each participant is then asked to pick which of the two versions they prefer. Each trace is repeated eight times, and the repetitions across traces are also randomized. The entire experiment lasts about one hour for each participant.
@@ 7. Evaluation
We first show that the subjective rendering quality of our method is statistically no-worse than Mini-Splatting-D (Sec. 7.1). We then show that we provide a better speed-quality trade-off than virtually all baselines on a mobile GPU (Sec. 7.2); hardware acceleration improves the speed even further (Sec. 7.3). We out-perform other FR methods (Sec. 7.4) and a prior PBNR accelerator (Sec. 7.5).
Fig. 13. Performance and objective rendering quality (PSNR, SSIM, LPIPS) comparison across the seven baselines and the three MetaSapiens variants on the mobile Volta GPU. 3DGS, Mini-Splatting-D, Mip-Splatting, and StopThePop are dense models, and the other three baselines are pruned models.
@@
7.2. GPU Results
We now show the performance results on the mobile Volta GPU on Nvidia Jetson AGX Xavier (xav, 2018), a representative mobile device for use-cases such as VR. Fig. 13 shows the results. We execute each model five times for each camera pose in each scene in all the datasets, and report the average FPS.
To put the performance results in context, we compare our variants with the baselines using three objective metrics, PSNR, SSIM, and LPIPS, which provide good quality measures for the region under the user’s gaze and are commonly reported in prior work. Using the objective metrics also allows us to scale up the study to more traces.
Our three variants provide better speed-quality trade-offs in virtually all metrics. Our slowest variant MetaSapiens-H is 1.9$\times$ faster than the fastest baseline while having better or similar objective quality. The fastest variant MetaSapiens-L is 7.9$\times$ faster than 3DGS, and can be up to 19.8$\times$ on the largest bicycle trace.
@@ 7.2. GPU Results
We now ablate the contribution of various performance-enhancing techniques. Fig. 12 shows the FPS (left $y$-axis) and PSNR (right $y$-axis) under: 1) the dense Mini-Splatting-D model, 2) MetaSapiens with only scale decay (SD; Sec. 3.3), 3) MetaSapiens with SD and CE-based pruning (Sec. 3.2), and 4) MetaSapiens with SD, CE, and FR (Sec. 4). We use the MetaSapiens-H model and obtain the FPS/PSNR results on Xavier averaged over all traces.
@@
Fig. 14. Speed-ups of different accelerator variants over the GPU baseline; each marker is a dataset trace.
Fig. 15. Speedup and area comparison between our hardware and GSCore under different configurations.
7.3. Results with Hardware Support
Speedup.
Our hardware support provides further performance improvements. Fig. 15 shows the speedups over GPU of: 1) the base accelerator (the uncolored in Fig. 8), 2) the accelerator with only Tile Merging, and 3) one with both Tile Merging and Incremental Pipelining. Each marker represents one of the 13 dataset traces. We use the MetaSapiens-H model for evaluation. Overall, even the base accelerator achieves a 18.5$\times$ speedup (geomean), up to 24.8$\times$, compared to the GPU baseline across different datasets.
The introduction of Tile Merging consistently improves performance, because tile merging mitigates the load imbalance across tiles. On top of that, Incremental Pipelining completely addresses the load imbalance in FR. Overall, MetaSapiens-TM-IP combines both techniques and achieves an average 20.9$\times$ (up to 27.7$\times$) speedup.
Energy Savings.
We summarize the energy results. Our base accelerator achieves a 54.4$\times$ energy reduction compared to the GPU baseline. MetaSapiens-TM-IP improves the energy saving to 56.8$\times$; this is primarily because with incremental pipelining we can afford to smaller SRAMs as line buffers, which reduces the energy consumption of SRAMs.
@@ 7.3. Results with Hardware Support
MetaSapiens-H 102.2 (0.81$\times$) 171.8 (1.06$\times$) 2.12 2.10 2.09 2.08
@@ 7.4. Comparison with Other FR Methods
MetaSapiens also out-performs the two FR baselines. Tbl. 1 compares the FPS (on the mobile Volta GPU), storage requirement, and the HVSQ metric across different quality regions/layers. The results are averaged across all the datasets.
@@
Our method selectively multi-versioned four trainable parameters out of about 60 (Sec. 4.2 and D in Fig. 7), so the additional storage requirement is small (about 6%). Note that MetaSapiens-H already reduces the dense model size to 16% as shown in Sec. 6.
@@
7.5. Comparison with GSCore
Our accelerator is based on GSCore (Lee et al., 2024a), but has a larger area (Sec. 6). This is because our baseline hardware (uncolored in Fig. 8) has 4$\times$ more Volume Rendering Cores compared to that of GSCore with 2$\times$ fewer sorting unit to balance the latency of different stages.
Fig. 15 compares the speedup over GPU and area between our architecture (with TM and IP) and GSCore; both hardware execute the MetaSapiens-H model on the flowers scene. We proportionally scale both GSCore and ours based on their own resource ratio. We consistently achieve higher speedups with a slightly smaller area against GSCore. For example, at an area of around 6 mm², MetaSapiens out-performs GSCore by 1.6$\times$. The higher performance of MetaSapiens is from the effectiveness of TM and IP that avoid stalling when pipelining over tiles. Our performance gain is more significant as the area increases, because the workload imbalance is more pronounced with a large amount of idle resource.
@@ 8. Related Work
The graphics community has long exploited FR for real-time rendering (Patney et al., 2016; Guenter et al., 2012; Chen et al., 2022a; Konrad et al., 2020; Kaplanyan et al., 2019; Krajancich et al., 2020; Chakravarthula et al., 2021; Sun et al., 2017; Singh et al., 2023). Particularly relevant to our work, researchers have started applying FR to neural rendering (Deng et al., 2022; Rolff et al., 2023b, a), such as Fov-NeRF (Deng et al., 2022). Our work differs from them in two key ways. First, these methods exclusively focus on NeRF, whereas we focus on PBNR, which is shown to be fundamentally more efficient that NeRF. Second, some (Deng et al., 2022) use the multi-model approach, similar to our MMFR baseline, which we out-perform (Sec. 7.4). Our FR method uses subsetting with selectively multi-versioning, addressing the performance overhead of evaluating multiple models (Sec. 4.1).
While conventional FR uses heuristics (e.g., blurring) to guide quality relaxation, recently researchers have investigated more principled ways to model human perception for quality relaxation (Freeman and Simoncelli, 2011; Walton et al., 2021). This work leverages such theoretical work and integrates it into the training framework to demonstrate its practical utility.
@@
Almost all existing work optimizing PBNR focuses on pruning, based on the observation that a considerable amount of points can be pruned without impacting the rendering quality. They usually do so by, e.g., explicitly training a mask to remove points (Lee et al., 2024b) or sorting points by their numerical contribution to pixel colors followed by removing low-contributing points (Fan et al., 2024; Girish et al., 2024; Niemeyer et al., 2024; Fang and Wang, 2024). People have also investigated non-pruning methods, such as vector quantization (Fan et al., 2024) and distillation (Lee et al., 2024b) techniques, to compress PBNR models.
@@
Neural Rendering Accelerators.
Significant work has been done on accelerating neural rendering (Lee et al., 2023; Rao et al., 2022; Li et al., 2023; Mubarik et al., 2023; Li et al., 2022; Fu et al., 2023; Feng et al., 2024), but most of them focuses on NeRF, a variant of neural rendering that PBNR aims to out-perform. GSCore (Lee et al., 2024a) accelerates 3DGS, a particular PBNR model.
Our work differs from prior PBNR accelerators in two ways. First, our hardware supports FR with minimal hardware augmentations. Second, we identify and address the load imbalance issue in PBNR, which is exacerbated by FR.